OpenStudy (anonymous):

Factorise the expression 5x^3+9x^2-96x-80 into three linear factories ,given (x+5) is a factor.

OpenStudy (anonymous):

Divide the given polynomial with (x+5), u will get the quotient. factorise the quotient, and u get the other 2 factors.

OpenStudy (anonymous):

how complete please ?

OpenStudy (anonymous):

Wait.

OpenStudy (ash2326):

We have \[5x^3+9x^2-96x-80\] It's given that x+5 is a factor, either we can divide the expression by x+5 to find a quadratic equation which will give us the two other factors or we could create x+5 factor in the expression and take it out to find our quadratic. I'll show you the second one.

sam (.sam.):

5x^(3)+9x^(2)-96x-80 (x-4)(x+5)(x+(4)/(5))

OpenStudy (dumbcow):

use synthetic division -5 | 5 9 -96 -80 -25 80 80 ----------------- 5 -16 -16 0 --> 5x^2 -16x -16

OpenStudy (ash2326):

\[5x^3+9x^2-96x-80\] Let's create x+5 we can write first term as \[5x^3=5x^2(x+5)-25x^2\] so \[5x^2(x+5)-25x^2+9x^2-96x-80\] or \[5x^2(x+5)-16x^2-96x-80\] Writing second term as \[-16x^2=-16x(x+5)+80x \] now we have \[5x^2(x+5)-16x(x+5)+80x-96x-80\] or \[5x^2(x+5)-16x(x+5)-16x-80=>5x^2(x+5)-16x(x+5)-16(x+5)\] so we get \[(x+5)(5x^2-16x-16)\] now let's factor \[(5x^2-16x-16\] we need to find factors of -16*5=-80 such that their sum is -16 -20 and 4 satisfy this so \[(5x^2-20x+4x-16=>5x(x-4)+4(x-4)\] we get \[(x-4)(5x+4)\] so (x-4) and (5x+4) are the two other factors

OpenStudy (anonymous):

you can try by using long division.. or try any value.. subtitube any number in x the it will be the factor of that equation ..

OpenStudy (anonymous):

thank you guys